A carpenter and a helper can do a certain job in 24 days. If the carpenter works 1.50 times as fast as the helper, how long would it take each, working alone, to do the job.
Answer by mananth(16949)

let helper take x hours to do the job alone
the carpenter will take x/(1.5 ) hours to do the job alone
x/(1.5) = 2x/3
In one day they both do
3/2x + 1/x of the job
3/2x + 1/x = 1/24
(3+2)/2x = 1/24
5*24= 2x
120=2x
x=60
helper does alone in 60 days
Carpenter does in 60/1.5 = 40 days